# Stanford CA 94305, USA

### v1

@@ 1 Introduction
We will look at several approaches to UQ that all require some extra knowledge. One approach is to suppose that $f$ belongs to a given function class ${\mathcal{F}}$, in which we can bound $\varepsilon_{n}=\varepsilon_{n}(f)$. Another is to use a model with random $\hat{\mu}_{n}$ and from there estimate the distribution of $\varepsilon_{n}$. In many cases we have asymptotic error estimation that becomes more precise as the amount of computation becomes infinite but is only approximate for finite computation.
@@
Since this article is to appear in an MCQMC proceedings, we assume some background knowledge that would take too much space to include here. For readers new to those topics, here are some bibliographic references that also include some history. Those topics include discrepancy [11], the ANOVA decomposition [69, Appendix A], weighted Hilbert spaces [22], digital QMC constructions [23] (such as Sobol’ sequences and Faure sequences) integration lattices [78] and their randomizations [49].
The contents of this paper are as follows. Section 2 introduces some notation along with basic strategies to get an estimate $\hat{\mu}$ of $\mu$. Section 3 describes the basics of uncertainty quantification. There are certificates, confidence intervals, and asymptotic confidence intervals. The results making the more desirable claims are harder to apply. Section 4 gives some classical certificates (bracketing inequalities) for numerical integration of convex functions along with a new one for completely monotone integrands using QMC points with nonnegative and nonpositive local discrepancy. Section 5 covers RQMC. It emphasizes asymptotic statistical confidence intervals as $n\to\infty$ and how hard they are to apply to RQMC with a small number $R$ of independent replicates of an RQMC rule on $n$ points. An empirical investigation in [52] showed that classical $t$-test intervals based on $R$ independent replicates are surprisingly reliable and there are now a few theoretical explanations to show why. Section 6 looks at a proposal by Tony Warnock and John Halton that in the notation of this paper is like using $R$ quasi-random replicates of a QMC rule instead of $R$ random replicates. Their quasi-standard error has had some empirical success and also has some failings. The idea might benefit from a further investigation. Section 7 describes the guaranteed automatic integration library which aims to sample adaptively until a specified error criterion $|\hat{\mu}-\mu|\leqslant\epsilon$ is met, either absolutely or probabilistically. Section 8 presents areas of current active interest in UQ for QMC: unbiased MCMC, normalizing flows and median of means. Conclusions in Section 9 describe how the present understanding of the problem has many gaps. It appears that there will be a role for the analysis of distributions that become symmetric as $n\to\infty$ without becoming Gaussian.
@@ 2 Notation and basic estimates of $\mu$
This paper surveys many different results from multiple literatures. As a result, it is necessary for some symbols to be used for different things. The specific use cases are different enough to be clear from context.
@@
then ${\mathbb{E}}((\hat{\mu}_{n}-\mu)^{2})=\sigma^{2}/n$ and the root mean squared error (RMSE) is $\sigma/\sqrt{n}$. The $n^{-1/2}$ rate is not as good as the ones for smooth low dimensional functions in [17] but remarkably it requires no differentiability and is the same in any dimension.
@@
is the local discrepancy at ${\bm{a}}$. The QMC counterpart to the law of large numbers is that if $f$ is Riemann integrable and $D_{n}^{*}\to 0$ as $n\to\infty$, then $\hat{\mu}_{n}\to\mu$. See [63] which also mentions a converse that if $\hat{\mu}_{n}\to\mu$ whenever $D_{n}^{*}\to 0$, then $f$ must be Riemann integrable. We will see below that if $f$ is of bounded variation in the sense of Hardy and Krause (written $f\in\mathrm{BVHK}$) then $\varepsilon_{n}=O(D_{n}^{*})$. Since it is possible to find points with $D_{n}^{*}=O(n^{-1+\epsilon})$ for any $\epsilon>0$, we can get asymptotically lower errors from QMC than the RMSE of MC.
@@ 3 Uncertainty quantification
Equation (2) falls short of being an uncertainty quantification because we ordinarily do not know either of the factors in it. It is expensive to compute $D_{n}^{*}$. The commonly used algorithm of [25] costs $O(n^{1+d/2})$. In the time it would take us to compute $D_{n}^{*}$ for QMC we could use $N=O(n^{1+d/2})$ points in MC. Then MC would get an RMSE of $O(n^{-1/2-d/4})$ in the time that QMC using known $D_{n}^{*}$ would get an error of $O(n^{-1+\epsilon})$. We could in principle precompute $D_{n}^{*}$ for a number of point sets, but then we would face the second problem: $V_{\mathrm{HK}}(f)$ is extremely hard to compute, far worse than $\mu$ itself. For smooth enough $f$, it is a sum of $12^{d}-$ 1 integrals of the absolute values of some mixed partial derivatives of $f$ [67]. As a result, (2) serves to show us that QMC can be much better than MC, but it does not generally let us verify that that has happened for given $f$ and $n$.
More modern bounds are derived for reproducing kernel Hilbert spaces (RKHSs) of integrands. The unanchored Sobolev spaces described in [22] are a prominent example. For every $u\subseteq 1{:}d$, define a weight $\gamma_{u}>0$ and then let ${\bm{\gamma}}$ represent all $2^{d}$ of those weights. The norm for this space is
@@
for weak partial derivatives $\partial^{u}f$. The error $|\hat{\mu}-\mu|$ is upper bounded by $\|f\|_{{\bm{\gamma}}}$ times a corresponding $L_{2}$ discrepancy measure defined by a reproducing kernel [22]. Product weights have $\gamma_{u}=\prod_{j\in u}\gamma_{j}$ for $\gamma_{j}>0$. When the $\gamma_{j}$ decay rapidly with increasing $j$, for example with $\gamma_{j}=j^{-\eta}$ for $\eta>1$ then strong tractability (see [79]) holds, in which the number of function evaluations required to get an error below $\epsilon\|f\|_{{\bm{\gamma}}}$ does not depend on the dimension $d$. For $\eta>2$, strong tractability holds with errors $O(n^{-1+\delta})$ for any $\delta>0$. This tractability property has been widely used to design better QMC point sets. However evaluating $\|f\|_{{\bm{\gamma}}}$ for an integrand $f$ is usually quite difficult and so, just like the Koksma-Hlawka inequality, the error bounds in weighted spaces do not provide an uncertainty quantification.
@@
We would prefer a non-asymptotic and nonparametric confidence interval. That would provide exactly $1-\alpha$ coverage probability without requiring that $f({\bm{x}})$ have a distribution belonging to a finite dimensional family, such as the Gaussian distributions. Unfortunately, exact nonparametric confidence intervals do not exist under conditions given in Bahadur and Savage [5]. They consider a set ${\mathcal{F}}$ of distributions on ${\mathbb{R}}$. Letting $\mu(F)$ be ${\mathbb{E}}(Y)$ when $Y\sim F\in{\mathbb{R}}$, their conditions are
@@
Then their Corollary 2 shows that a Borel set constructed based on $x_{1},\dots,x_{N}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$ that contains $\mu(F)$ with probability at least $1-\alpha$ also contains any other $m\in{\mathbb{R}}$ with probability at least $1-\alpha$. More precisely: we can get a confidence set, but not a useful one. They allow $N$ to be random so long as $\Pr(N<\infty)=1$.
@@
Suppose that independent $f({\bm{x}}_{i})$ satisfies known finite upper and lower bounds. Hoeffding’s inequality [46] then provides confidence intervals for $\sum_{i=1}^{n}f({\bm{x}}_{i})$ in terms of those bounds. In the MC context it is natural to assume that the bounds are the same for all $i=1,\dots,n$. Because the bounds are known we can make a linear adjustment to $f$ and then without loss of generality we may then suppose that $0\leqslant f({\bm{x}})\leqslant 1$ for all ${\bm{x}}\in[0,1]^{d}$. Then for $t>0$, Hoeffding’s inequality gives
@@
This holds for any distribution of $f({\bm{x}})$ with $\Pr(0\leqslant f({\bm{x}})\leqslant 1)=1$. It is possible because bounded random variables do not satisfy Bahadur and Savage’s condition (ii).
@@ 4 Certificates and bracketing
A certificate is a (finite) computable number that is mathematically guaranteed to be no smaller than $\varepsilon_{n}$. This concept is widely used in convex optimization [7]. When the goal is to minimize a convex function $f$ over ${\bm{x}}\in C$ for a convex set $C$, we can be sure that the minimum is no larger than $\min(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{n}))$ once $f$ has been evaluated at points ${\bm{x}}_{i}$ for $i=1,\dots,n$. Because $f$ does not go below any of its supporting hyperplanes the minimum must not be below the function value where any set of those hyperplanes intersect. That can provide a computable lower bound for the minimum of $f$. The user then has mathematical certainty that the answer lies between two computable numbers $a$ and $b$. If we take $(a+b)/2$ as the estimate then we know that the error is no larger than $(b-a)/2$.
@@
The situation is very different for multivariate problems. Then there is a curse of dimension from [6]. Here is a sketch based on [24]. Suppose that the partial derivative of $f$ taken $a_{j}\geqslant 0$ times with respect to component $x_{j}$ is never larger than some $M\in(0,\infty)$ for any ${\bm{x}}\in[0,1]^{d}$ when $\sum_{j=1}^{d}a_{j}\leqslant r$. Then there still exists $B>0$ such that for any rule of the form $\sum_{i=1}^{n}w_{i}f({\bm{x}}_{i})$ for $w_{i}\in{\mathbb{R}}$ and ${\bm{x}}_{i}\in[0,1]^{d}$ there is $f$ with $|\hat{\mu}(f)-\mu(f)|>Bn^{-r/d}$. For randomized points ${\bm{x}}_{i}$, the RMSE cannot be below $B^{\prime}n^{-1/2-r/d}$ for some $B^{\prime}>0$. For large $d$, we cannot simply ignore the integration error. It is also not easy to use methods tuned for $r\geqslant 3$. Dimov [24] describes them as ‘sophisticated’. With lower bounds that bad, upper bounds and estimates of upper bounds can only be worse for those function classes.
@@
This method can be applied to domains that are partitioned into simplices, allowing us to generalize the rule in (12) while reusing any function evaluations that are on a vertex of more than one simplex. For integration over $[0,1]^{d}$ any such rule would require at least the $2^{d}$ function evaluations on the corners of the unit cube, and would therefore not scale well to large dimensions.
@@
holds for every nonempty $u\subseteq 1{:}d$ and all ${\bm{x}},{\bm{z}}\in[0,1]^{d}$ with ${\bm{x}}\leqslant{\bm{z}}$ componentwise. Taking $u=\{j\}$ this requires that $f({\bm{x}}_{-j}{:}{\bm{z}}_{j})\geqslant f({\bm{x}})$ so $f$ is nondecreasing in each of the $d$ variables. For $u=\{j,k\}$ for $j\neq k$ we see that $f({\bm{x}}_{-j}{:}{\bm{z}}_{j})-f({\bm{x}})$ is nondecreasing in variable $k$. Generally an $r$-fold difference of differences is nonnegative.
@@
If ${\bm{x}}_{1},\dots,{\bm{x}}_{n}$ have NPLD and either all ${\bm{x}}_{i}\in[0,1)^{d}\cup\{{\bm{1}}]\}$, or $\nu$ is absolutely continuous with respect to Lebesgue measure, then
@@
While the $O(n^{-2/d+\epsilon})$ rate shows a strong dimension effect, it is better than the rate for bounded functions that are simply nondecreasing in each variable individually. Deterministic methods cannot be better than $O(n^{-1/d})$, and random methods cannot have a better RMSE than $O(n^{-1/d-1/2})$. See [75] for precise statements with proof.
@@ 5 Randomized quasi-Monte Carlo
The empirical findings motivated subsequent work in [73] which shows that for random linear scrambling of a digital net in base $2$ with a digital shift, the skewness of $\hat{\mu}_{n}$ is $\gamma_{n}=O(n^{\epsilon})$ for any $\epsilon>0$, so it is almost $O(1)$. Furthermore, under a model with randomly chosen generator matrices the skewness is $O(n^{-1/2+\epsilon})$. The kurtosis of $\hat{\mu}$ from random linear scrambling of a digital net in base $2$ is known to diverge to $\infty$ as $n\to\infty$ for an analytic function on $[0,1]$ by a finding in [72, Section 3]. The reasoning is as follows. There is an event of probability $\Omega(1/n)$ that gives a squared error of $\Omega(1/n^{2})$ (for smooth $f^{\prime}$ that does not integrate to zero). That same event on its own contributes $\Omega(n^{-5})$ to the fourth power of the error. The expected squared error is $O(n^{-3+\epsilon})$ and so $\kappa_{n}+3=\Omega(n^{-5})/O(n^{-6+2\epsilon})=\Omega(n^{1-2\epsilon})$. An integrand on $[0,1]^{d}$ for $d>1$ with a smooth contribution from a one dimensional main effect will have the same diverging kurtosis. Having $\kappa_{n}\to\infty$ completely rules out a CLT for $\hat{\mu}_{n}$ from matrix scrambling with a digital shift.

### v2

@@ 1 Introduction
We will look at several approaches to UQ that all require some extra knowledge. One approach is to suppose that $f$ belongs to a given function class ${\mathcal{F}}$, in which we can bound $\varepsilon_{n}=\varepsilon_{n}(f)$. Another is to use a model with random $\hat{\mu}_{n}$ where we can get bounds on the distribution of $\varepsilon_{n}$. Sometimes we can only get asymptotically justified bounds as the amount of computation grows to infinity.
@@
Since this article is to appear in an MCQMC proceedings, we assume some background knowledge that would take too much space to include here. For readers new to those topics, here are some bibliographic references that also include some history. Those topics include discrepancy [11], the ANOVA decomposition [69, Appendix A], weighted Hilbert spaces [22], digital QMC constructions [23] (such as Sobol’ sequences and Faure sequences) integration lattices [79] and their randomizations [49]. Some further references are given in context.
The contents of this paper are as follows. Section 2 introduces some notation along with basic strategies to get an estimate $\hat{\mu}$ of $\mu$. Section 3 describes the basics of uncertainty quantification. There are certificates, confidence intervals, and asymptotic confidence intervals. The results making the more desirable claims are harder to apply. Section 4 gives some classical certificates (bracketing inequalities) for numerical integration of convex functions along with a new one for completely monotone integrands using QMC points with nonnegative and nonpositive local discrepancy. Section 5 covers RQMC. It emphasizes asymptotic statistical confidence intervals as $n\to\infty$ and how hard they are to apply to RQMC with a small number $R$ of independent replicates of an RQMC rule on $n$ points. An empirical investigation in [52] showed that classical $t$-test intervals based on $R$ independent replicates are surprisingly reliable and there are now a few theoretical explanations to show why. Section 6 looks at a proposal by Tony Warnock and John Halton that in the notation of this paper is like using $R$ quasi-random replicates of a QMC rule instead of $R$ random replicates. Their quasi-standard error has had some empirical success and also has some failings. The idea might benefit from a further investigation. Section 7 describes the guaranteed automatic integration library which aims to sample adaptively until a specified error criterion $|\hat{\mu}-\mu|\leqslant\epsilon$ is met, either absolutely or probabilistically. Section 8 presents areas of current active interest in UQ for QMC: unbiased MCMC, normalizing flows, median of means and results about $R$ growing with $n$. Conclusions in Section 9 describe how the present understanding of the problem has many gaps. It appears that there will be a role for the analysis of distributions that become symmetric as $n\to\infty$ without becoming Gaussian.
@@ 2 Notation and basic estimates of $\mu$
This paper surveys many different results from multiple literatures. As a result, it is necessary for some symbols to be used with more than one meaning. The specific use cases are different enough to be clear from context.
@@
then ${\mathbb{E}}((\hat{\mu}_{n}-\mu)^{2})=\sigma^{2}/n$ and the root mean squared error (RMSE) is $\sigma/\sqrt{n}$. The $n^{-1/2}$ rate is not as good as the ones for smooth low dimensional functions in [17] but remarkably, it requires no differentiability and is the same in any dimension.
@@
is the local discrepancy at ${\bm{a}}$. The QMC counterpart to the law of large numbers is that if $f$ is Riemann integrable and $D_{n}^{*}\to 0$ as $n\to\infty$, then $\hat{\mu}_{n}\to\mu$. See [63] which also mentions a converse that if $\hat{\mu}_{n}\to\mu$ whenever $D_{n}^{*}\to 0$, then $f$ must be Riemann integrable. We will see below that if $f$ is of bounded variation in the sense of Hardy and Krause (written $f\in\mathrm{BVHK}$) then $\varepsilon_{n}=O(D_{n}^{*})$. Since it is possible to find points with $D_{n}^{*}=O(n^{-1+\epsilon})$ for any $\epsilon>0$, we can get asymptotically lower errors from QMC than the RMSE of MC. The $\epsilon$ hides powers of $\log(n)$. Those are present in some worst cases, but they do not seem to appear in applications [70].
@@ 3 Uncertainty quantification
Equation (2) falls short of providing an uncertainty quantification because we ordinarily do not know either of the factors in it. It is expensive to compute $D_{n}^{*}$. The commonly used algorithm of [25] costs $O(n^{1+d/2})$. In the time it would take us to compute $D_{n}^{*}$ for QMC we could get $N=\Omega(n^{1+d/2})$ points in MC. Then MC would get an RMSE of $O(n^{-1/2-d/4})$ in the time that QMC using known $D_{n}^{*}$ would get an error of $O(n^{-1+\epsilon})$. We could in principle precompute $D_{n}^{*}$ for a number of point sets, but then we would face the second problem: $V_{\mathrm{HK}}(f)$ is extremely hard to compute, far worse than $\mu$ itself. For smooth enough $f$, it is a sum of $12^{d}-$ 1 integrals of the absolute values of some mixed partial derivatives of $f$ [67]. For less smooth $f$, the problem is even harder. As a result, (2) serves to show us that QMC can be much better than MC, but it does not generally let us verify that that has happened for given $f$ and $n$.
More modern bounds are derived for reproducing kernel Hilbert spaces (RKHSs) of integrands. The unanchored Sobolev spaces described in [22] are a prominent example. For every $u\subseteq 1{:}d$, define a weight $\gamma_{u}>0$ and then let ${\bm{\gamma}}$ represent all $2^{d}$ of those weights. The squared norm in this space is
@@
for weak partial derivatives $\partial^{u}f$. The error $|\hat{\mu}-\mu|$ is upper bounded by $\|f\|_{{\bm{\gamma}}}$ times a corresponding $L_{2}$ discrepancy measure defined by a reproducing kernel [22]. The widely used product weights have $\gamma_{u}=\prod_{j\in u}\gamma_{j}$ for $\gamma_{j}>0$. When $\gamma_{j}$ decays rapidly with increasing $j$, for example with $\gamma_{j}=j^{-\eta}$ for $\eta>1$, then strong tractability (see [80]) holds, in which the number of function evaluations required to get an error below $\epsilon\|f\|_{{\bm{\gamma}}}$ does not depend on the dimension $d$. For $\eta>2$, strong tractability holds with errors $O(n^{-1+\delta})$ for any $\delta>0$. This tractability property has been widely used to design better QMC point sets. However evaluating $\|f\|_{{\bm{\gamma}}}$ for an integrand $f$ is usually quite difficult and so, just like the Koksma-Hlawka inequality, the error bounds in weighted spaces do not provide an uncertainty quantification.
@@
We would prefer a non-asymptotic and nonparametric confidence interval. That would provide exactly $1-\alpha$ coverage probability without requiring that $f({\bm{x}})$ have a distribution belonging to a finite dimensional family, such as the Gaussian distributions. Unfortunately, exact nonparametric confidence intervals do not exist under conditions given in Bahadur and Savage [5]. They consider a set ${\mathcal{F}}$ of distributions on ${\mathbb{R}}$. Letting $\mu(F)$ be ${\mathbb{E}}(Y)$ when $Y\sim F\in{\mathbb{R}}$, their conditions are:
@@
Then their Corollary 2 shows that a Borel set constructed based on $Y_{1},\dots,Y_{N}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$ that contains $\mu(F)$ with probability at least $1-\alpha$ also contains any other $m\in{\mathbb{R}}$ with probability at least $1-\alpha$. More precisely: we can get a confidence set, but not a useful one. They allow $N$ to be random so long as $\Pr(N<\infty)=1$.
@@
Suppose that independent $f({\bm{x}}_{i})$ satisfies known finite upper and lower bounds. Hoeffding’s inequality [46] then provides confidence intervals for $(1/n)\sum_{i=1}^{n}f({\bm{x}}_{i})$ in terms of those bounds. In the MC context it is natural to assume that the bounds are the same for all $i=1,\dots,n$. Because the bounds are known we can make a linear adjustment to $f$ and then without loss of generality we may then suppose that $0\leqslant f({\bm{x}})\leqslant 1$ for all ${\bm{x}}\in[0,1]^{d}$. Then for $t>0$, Hoeffding’s inequality gives
@@
This holds for any distribution of $f({\bm{x}})$ with $\Pr(0\leqslant f({\bm{x}})\leqslant 1)=1$. It is possible because that set of bounded random variables does not satisfy Bahadur and Savage’s condition (ii).
A very natural setting with known bounds arises when $f({\bm{x}})$ takes the value $1$ for ${\bm{x}}\in A$ and 00 otherwise. Then $\mu$ can be interpreted as the probability that ${\bm{x}}\in A$ when ${\bm{x}}\sim{\mathbb{U}}[0,1]^{d}$.
@@ 4 Certificates and bracketing
A certificate is a (finite) computable number that is mathematically guaranteed to be no smaller than $\varepsilon_{n}$. This concept is widely used in convex optimization [7]. When the goal is to minimize a convex function $f$ over ${\bm{x}}\in C$ for a convex set $C$, we can be sure that the minimum is no larger than $\min(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{n}))$ which becomes known once $f$ has been evaluated at points ${\bm{x}}_{i}$ for $i=1,\dots,n$. Because $f$ does not go below any of its supporting hyperplanes the minimum must not be below the function value where any set of those hyperplanes intersect. That can provide a computable lower bound for the minimum of $f$. The user then has certainty that the answer lies between two computable numbers $a$ and $b$. If we take $(a+b)/2$ as the estimate then we know that the error is no larger than $(b-a)/2$.
@@
The situation is very different for multivariate problems. Then there is a curse of dimension from [6]. Here is a sketch based on [24]. Suppose that the partial derivative of $f$ taken $a_{j}\geqslant 0$ times with respect to component $x_{j}$ is never larger than some $M\in(0,\infty)$ for any ${\bm{x}}\in[0,1]^{d}$ when $\sum_{j=1}^{d}a_{j}\leqslant r$. Then there still exists $B>0$ such that for any rule of the form $\sum_{i=1}^{n}w_{i}f({\bm{x}}_{i})$ for $w_{i}\in{\mathbb{R}}$ and ${\bm{x}}_{i}\in[0,1]^{d}$ there is $f$ with $|\hat{\mu}(f)-\mu(f)|>Bn^{-r/d}$. For randomized points ${\bm{x}}_{i}$, the RMSE cannot be below $B^{\prime}n^{-1/2-r/d}$ for some $B^{\prime}>0$. For large $d$, we cannot simply ignore the integration error. It is also not easy to use methods tuned for $r\geqslant 3$. Dimov [24] describes them as ‘sophisticated’.
@@
This method can be applied to domains that are partitioned into simplices, allowing us to generalize the rule in (12) while reusing any function evaluations that are on a vertex of more than one simplex. For integration over $[0,1]^{d}$ partitioning this domain into simplices would require at least the $2^{d}$ function evaluations on the corners of the unit cube, and would therefore not scale well to large dimensions.
@@
holds for every nonempty $u\subseteq 1{:}d$ and all ${\bm{x}},{\bm{z}}\in[0,1]^{d}$ with ${\bm{x}}\leqslant{\bm{z}}$ componentwise. Taking $u=\{j\}$ this requires that $f({\bm{x}}_{-j}{:}{\bm{z}}_{j})\geqslant f({\bm{x}})$ so $f$ is nondecreasing in each of the $d$ variables. For $u=\{j,k\}$ for $j\neq k$ we see that $f({\bm{x}}_{-j}{:}{\bm{z}}_{j})-f({\bm{x}})$ is nondecreasing in $x_{k}$. Generally an $r$-fold difference of differences is nonnegative.
@@
If ${\bm{x}}_{1},\dots,{\bm{x}}_{n}$ have NPLD and either all ${\bm{x}}_{i}\in[0,1)^{d}\cup\{{\bm{1}}\}$, or $\nu$ is absolutely continuous with respect to Lebesgue measure, then
@@
While the $O(n^{-2/d+\epsilon})$ rate shows a strong dimension effect, it is better than the rate for deterministic sampling of bounded functions that are simply nondecreasing in each variable individually. Deterministic methods cannot be better than $O(n^{-1/d})$, and random methods cannot have a better RMSE than $O(n^{-1/d-1/2})$. See [76] for precise statements with proof.
@@ 5 Randomized quasi-Monte Carlo
The empirical findings motivated subsequent work in [74] which shows that for random linear scrambling of a digital net in base $2$ with a digital shift, the skewness of $\hat{\mu}_{n}$ is $\gamma_{n}=O(n^{\epsilon})$ for any $\epsilon>0$, so it is almost $O(1)$. Furthermore, under a model with randomly chosen generator matrices the skewness is $O(n^{-1/2+\epsilon})$. The kurtosis of $\hat{\mu}$ from random linear scrambling of a digital net in base $2$ is known to diverge to $\infty$ as $n\to\infty$ for an analytic function on $[0,1]$ by a finding in [73, Section 3]. The reasoning is as follows. There is an event of probability $\Omega(1/n)$ that gives a squared error of $\Omega(1/n^{2})$ (for smooth $f^{\prime}$ that does not integrate to zero). That same event on its own contributes $\Omega(n^{-5})$ to the fourth power of the error. The expected squared error is $O(n^{-3+\epsilon})$ and so $\kappa_{n}+3=\Omega(n^{-5})/O(n^{-6+2\epsilon})=\Omega(n^{1-2\epsilon})$. An integrand on $[0,1]^{d}$ for $d>1$ with a smooth contribution from a one dimensional main effect will have the same diverging kurtosis. Having $\kappa_{n}\to\infty$ completely rules out a CLT for $\hat{\mu}_{n}$ from matrix scrambling with a digital shift.